OpenStudy (anonymous):

30 basketballs (15 Spalding, 10 Wilson, and 5 other brand-name balls) are on a basketball court. Barry Wood closes his eyes and arbitrarily picks up a ball from the court. Determine the probability that the ball selected is a Wilson or a Spalding

OpenStudy (gw2011):

Probability of Wilson (W)=10/30 Probability of Spalding (S)=15/30 Probability of W or S=10/30 + 15/30 = 25/30 = 5/6

OpenStudy (anonymous):

Problem 3 frequency distribution Number of club members Hours per day 0-2 2 3-5 4 6-8 1 9-12 1 Second example of Frequency distribution Number of hours Club members 0-5 Mike, Ricardo, Sue, Said 5-10 Nicki, Juan, Chris,Alex tll u help me correct it p

OpenStudy (anonymous):

Problem 3 frequency distribution Number of club members Hours per day 0-2 2 3-5 4 6-8 1 9-12 1 Second example of Frequency distribution Number of hours Club members 0-5 Mike, Ricardo, Sue, Said 5-10 Nicki, Juan, Chris,Alex tll u help me correct it p

OpenStudy (anonymous):

Problem 3 frequency distribution Number of club members Hours per day 0-2 2 3-5 4 6-8 1 9-12 1 Second example of Frequency distribution Number of hours Club members 0-5 Mike, Ricardo, Sue, Said 5-10 Nicki, Juan, Chris,Alex tll u help me correct it p

OpenStudy (anonymous):

Problem 3 frequency distribution Number of club members Hours per day 0-2 2 3-5 4 6-8 1 9-12 1 Second example of Frequency distribution Number of hours Club members 0-5 Mike, Ricardo, Sue, Said 5-10 Nicki, Juan, Chris,Alex tll u help me correct it p

OpenStudy (anonymous):

Problem 3 frequency distribution Number of club members Hours per day 0-2 2 3-5 4 6-8 1 9-12 1 Second example of Frequency distribution Number of hours Club members 0-5 Mike, Ricardo, Sue, Said 5-10 Nicki, Juan, Chris,Alex tll u help me correct it p

OpenStudy (anonymous):

Problem 3 frequency distribution Number of club members Hours per day 0-2 2 3-5 4 6-8 1 9-12 1 Second example of Frequency distribution Number of hours Club members 0-5 Mike, Ricardo, Sue, Said 5-10 Nicki, Juan, Chris,Alex tll u help me correct it p

OpenStudy (anonymous):

Problem 3 frequency distribution Number of club members Hours per day 0-2 2 3-5 4 6-8 1 9-12 1 Second example of Frequency distribution Number of hours Club members 0-5 Mike, Ricardo, Sue, Said 5-10 Nicki, Juan, Chris,Alex tll u help me correct it p

OpenStudy (anonymous):

Problem 3 frequency distribution Number of club members Hours per day 0-2 2 3-5 4 6-8 1 9-12 1 Second example of Frequency distribution Number of hours Club members 0-5 Mike, Ricardo, Sue, Said 5-10 Nicki, Juan, Chris,Alex tll u help me correct it p

OpenStudy (anonymous):

Problem 3 frequency distribution Number of club members Hours per day 0-2 2 3-5 4 6-8 1 9-12 1 Second example of Frequency distribution Number of hours Club members 0-5 Mike, Ricardo, Sue, Said 5-10 Nicki, Juan, Chris,Alex tll u help me correct it p

OpenStudy (anonymous):

is either of these on the right track? I attached it so just look at the one that is for problem 3 and if u can the truth table too!

OpenStudy (anonymous):

is either of these on the right track? I attached it so just look at the one that is for problem 3 and if u can the truth table too!